OpenStudy (anonymous):

Evaluate -13 + |3.1 - 4.2|.

OpenStudy (anonymous):

-13 + |3.1 - 4.2| = -13 + | -1.1| = -13 + 1.1 = -11.9

OpenStudy (anonymous):

Absolute value cannot be negative and it is greater than or equal to zero and since 3.1<4.2 So you must multiply by -1 to remove absolute value sign. =-13+(-1(3.1-4.2)) And then @RolyPoly's answer :)